ACTIVITY: The average cholesterol content of a certain canned good is 215 mg, and the standard deviation is 15 mg. Assume that the variable is normally distributed. If a sample of 25 canned goods are selected, what is the probability that the mean of the sample will be larger than 220 mg


1
Expert's answer
2021-06-01T08:51:25-0400

Let "X=" the average cholesterol content of a certain can goods in milligrams.


"X\sim N(\mu, \sigma^2/n)\\\"

Then "Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0. 1)"


Now, Given

"X= 220\\\mu=215\\\sigma = 15\\n=25"


"z=\dfrac{X-\mu}{\sigma/\sqrt n}=\dfrac{220-215}{15/\sqrt{25}}=\dfrac{25}{15}=1.67"


So, probability that mean of sample is greater than 220 mg is

"P(Z>220)=P(z>1.67)"


So, from standard normal distribution table:

"P(Z>1.67)=0.0475"


Answer:

The probability that mean of sample is greater than 220 mg is 0.0475